# Generative AI for Immersive Communication: The Next Frontier in Internet-of-Senses Through 6G

### v1

@@ Abstract
Over the past two decades, the Internet-of-Things (IoT) has been a transformative concept, and as we approach 2030, a new paradigm known as the Internet of Senses (IoS) is emerging. Unlike conventional Virtual Reality (VR), IoS seeks to provide multi-sensory experiences, acknowledging that in our physical reality, our perception extends far beyond just sight and sound; it encompasses a range of senses. This article explores existing technologies driving immersive multi-sensory media, delving into their capabilities and potential applications. This exploration includes a comparative analysis between conventional immersive media streaming and a proposed use case that leverages semantic communication empowered by generative Artificial Intelligence (AI). The focal point of this analysis is the substantial reduction in bandwidth consumption by 99.93% in the proposed scheme. Through this comparison, we aim to underscore the practical applications of generative AI for immersive media while addressing the challenges and outlining future trajectories.
@@ I Introduction
The advent of the 5th generation (5G) mobile networks and the recent advancements in computing technologies have redefined the concept of internet from basic connectivity to a more advanced digital experience, transitioning from merely faster communication into an immersive interaction with the digital realm. This concept has been recently introduced under the umbrella of Metaverse and digital twins, and has opened up a wide range of applications including VR, Augmented Reality (AR), holoportation, and tele-operation, among others. Within this realm, four main underpinnings have been remarked as paradigms for linking the cyber and physical worlds, namely, connected intelligent machines, a digitized programmable world, connected sustainable world, and the IoS [1]. The IoS concept is set to revolutionize the digital interactions by creating a fully immersive environment that transcends traditional boundaries. By integrating sensory experiences such as sight, sound, touch, smell, and taste into the digital realm, this technology promises a more engaging cyber world, where virtual experiences are as rich and multi-dimensional as the physical world.
Figure 1: Key concepts of IoS
@@
Human beings experience the world through different senses, by perceiving sensory signals that are integrated or segregated in the brain. If these senses, especially haptic feedback, are accurately represented to be coherent with the real world, they can positively influence actions and behaviors, such as reaction time and detection [2]. Within this context, the IoS technology will allow individuals to experience a wide range of sensations remotely, revolutionizing various verticals, including industry, healthcare, networking, education, and tourism, to name a few. In order to reap the full potential of the IoS technology, numerous challenges need to be tackled to achieve a fully immersive multisensory experience. These challenges are pertinent to temporal synchronization of multiple media, addressing motion sickness, ensuring high throughput, and minimizing the End-to-End (E2E) latency. The collection of data from various sensor modalities, such as visual, audio, and haptic, plays a vital role in crafting a multisensory experience, in which this data can be synchronized at either the source or the destination (i.e., end devices or edge servers). The failure of virtual experiences to truly replicate our senses introduces confusion in human brains, leading to symptoms like nausea, dizziness, and migraines. To mitigate these drawbacks, it is crucial to enhance the realism of virtual sensations and reduce latency in VR/AR devices, thereby minimizing latency between different modalities and avoiding its mismatch [3]. Furthermore, for accurate control purposes over a distance of up to one mile and to prevent the occurrence of motion sickness, it is crucial to transmit the sensory information at extremely low E2E latency, ideally within 1-10 millisecond [4].
@@
With respect to the Key Performance Indicators for reliable communication of immersive media in IoS, it was demonstrated that future 6G networks should realize an E2E latency performance within the range of 1 ms for high-quality video streaming and haptic signals, with data rate requirements that go from tens Mbps to 1 Tbps and reliability performance of $10^{-7}$ [5]. Also, while taste and smell signals requirements are less stringent than videos and haptics, it is essential to realize a perfect synchronization among signals from different senses to achieve the full potential of the IoS. Among various technologies, semantic communication emerges as a promising candidate to achieve ultra-low latency communication through communicating the meanings/semantics of transmitted messages instead of communicating the whole signal, yielding faster and bandwidth efficient transmission.
@@
Accordingly, LLMs are envisioned to evolve into the cognitive hub of the IoS, addressing intricate challenges like synchronization and compression through the estimation from partial modalities and a communication enabled by semantic understanding. Additionally, LLMs are poised to enhance machine control intelligence, thereby improving reliability in teleoperations, through managing various data modalities pertinent to the user and environmental senses, as illustrated in Fig. 1.
@@
In this paper, we aim to set the scene for the integration of LLMs and the IoS technology, in which we develop a case study to demonstrate the benefits that can be obtained from exploiting the capabilities of LLMs in enhancing the latency performance of immersive media communication. In particular, we conceptualize the 360 video streaming from a Unmanned Aerial Vehicle (UAV) as a semantic communication task. Initially, we employ object detection and image-to-text captioning to extract semantic information (text) from the input 360 frame. Subsequently, this generated textual information is transmitted to the edge server. In the edge server, an LLM is utilized to produce code compatible with A-frame, facilitating the display of the corresponding image through Three-Dimensional (3D) virtual objects on the Head Mounted Device (HMD). Lastly, the generated code is sent to the receiver, allowing for the direct rendering of the 3D virtual content on the HMD.
The contributions of this paper are summarized as follows:
@@
  • Assess the quality of the generated 3D objects from our system compared to the captured 360 video images using reverse image-to-text, followed by text comparison through BERT model.
The remainder of this paper is organized as follows. Section II introduces the IoS and discusses its necessity. In Section III, an overview of the development of MLLMs and their applications to IoS is discussed. Section IV presents a case study with a proposed testbed, which is implemented and analyzed. Finally, Section VII highlights challenges and suggests directions for future research.
@@ II-A Immersive All-Sense Communication
Internet-of-Touch. Haptic sensation refers to the sense of touch, known as tactile sensation, and it enhances immersive multimedia by allowing individuals to feel physical sensations, such as interactions with objects and movements (kinesthetic sensation). In VR training or teleoperation, haptics replicate touch, which is crucial for tasks like surgery. Achieving optimal haptic experiences requires addressing minimal response times and low latency in synchronization with other sensed media, such as audio and video. Haptic interfaces employ various technologies to deliver tactile sensations, ranging from simple vibration feedback to more complex systems providing force feedback, pressure sensitivity, or even localized temperature changes. Devices like haptic gloves, exoskeletons, or tactile feedback controllers enable users to touch, grasp, and interact with virtual objects in a natural and intuitive manner.
Internet-of-Taste. Gustatory perception involves the intricate process of detecting and interpreting flavors. While traditional VR primarily focuses on visual and auditory stimuli, incorporating taste into the virtual environment has the potential to enhance sensory engagement, leading to more realistic and immersive experiences. The technology underlying gustatory interfaces centers on the controlled stimulation of taste receptors. Various approaches are being explored, such as electrically stimulating taste buds [9] or delivering taste-related chemical compounds directly to the mouth [10]. However, it is crucial to note that replicating the sense of taste is the most complex, as it closely depends on other sensations. Presently, the technology is still in the laboratory demonstration stage.
@@ III-A Advancement of Language Models
The progress in Natural Language Processing (NLP) research has led to the development of models such as GPT-2, BART [13], and BERT [14]. These models have sparked a new race to construct more efficient models with large-scale architectures, encompassing hundreds of billions of parameters. The most popular architecture is the decoder-only, including LLMs like GPT-3, Chinchilla and LaMDA. Following the release of open-source LLMs like OPT and BLOOM, more efficient open-source models have been recently introduced, such as Vicuna, Phi-1/2, LLaMa, FALCON, Mistral, and Mixtral111 https://huggingface.co/docs/transformers/model_doc/mixtral. This later follows the Mixture of Expects (MoE) architecture and training process initially proposed in MegaBlocks [15]. Despite having fewer parameters, these models fine-tuned on high-quality datasets, have demonstrated compelling performance on various NLP tasks, surpassing their larger counterparts. Furthermore, instruction tuning the foundation models on high-quality instruction datasets enables versatile capabilities like chat and code source generation, etc. The LLM have also shown unexpected capabilities of learning from the context (i.e., prompts), referred to as In-Context Learning (ICL).
@@ III-B Multimodal large language models
In this era of MLLMs, GPT-4 has demonstrated remarkable performance in vision-language tasks encompassing comprehension and generation. Nevertheless, in addition to its intricate nature, the technical details of GPT-4 remain undisclosed, and the source code is not publicly available, impeding direct modifications and enhancements [17]. To address these challenges, the MiniGPT-4 model was proposed. This model combines a vision encoder (ViT-G/14 and Q-Former) with the Vicuna LLM, utilizing only one projection layer to align visual features with the language model while keeping all other vision and language components frozen. The model is first trained on image-language datasets, then finetuned on high-quality image description pairs (3,500) to improve the naturalness of the generated language and its usability. The TinyGPT-V vision model, introduced by Yuan et al. [18], addresses computational complexity, necessitating only a 24GB GPU for training and an 8GB GPU for inference. The architecture of TinyGPT-V closely resembles that of MiniGPT-4, incorporating a novel linear projection layer designed to align visual features with the Phi-2 language model, which boasts only 2.7 billion parameters. The TinyGPT-V model undergoes a sophisticated training and fine-tuning process in four stages, where both the weights of the linear projection layers and the normalization layers of the language model are updated. As the process progresses, instruction datasets are incorporated in the third stage, and multi-task learning is employed during the fourth stage.
The second step in developing LLMs is fine-tuning the model on instruction datasets to teach models to better understand human intentions and generate accurate responses. The InstructBLIP [19] is built through instruct tuning of the pre-trained BLIP-2 model on 26 instruction datasets grouped into 11 tasks. During the instruction tuning process, the LLM and the image encoder are maintained frozen, while only the Q-former undergoes fine-tuning. Furthermore, the instructions are input to both the frozen LLM and the Q-Former. Notably, InstructBLIP exhibits exceptional performance across various vision-language tasks, showcasing remarkable generalization capabilities on unseen data. Moreover, when employed as the model initialization for individual downstream tasks, InstructBLIP models achieve state-of-the-art fine-tuning performance. InstructionGPT-4 [20] is a vision model fine-tuned on a small dataset comprising only 200 examples, which represents approximately 6% of the instruction-following data used in the alignment dataset for MiniGPT-4. The study highlights that fine-tuning the vision model on a high-quality instruction dataset enables the generation of superior output compared to MiniGPT-4.
@@
IV Case Study
IV-A Use case description
@@
To comprehend the challenges at hand and explore potential solutions, let us immerse ourselves in the following scenario. Consider John, a surveillance teleoperator, confronted with immense challenges as he endeavors to remotely operate drones with a First-Person View (FPV). His task involves teleoperating the UAV beyond the visual line of sight, navigating through a dense forest using Beyond 5G (B5G) networks. Navigating the forest terrain through FPV proves to be incredibly challenging, primarily due to two significant reasons.
  • On the one hand, the environment is characterized by bandwidth limitations that degrade the video stream perceived by John through FPV. This deterioration can significantly impact the teleoperation of the UAV and may lead to potentially catastrophic consequences.
  • On the other hand, the UAV is equipped only with a 360 camera, limiting John to receiving video and sound solely through the FPV. However, this does not entirely convey the environment in which the UAV is flying. Achieving complete immersion in controlling the UAV in such a challenging environment would necessitate additional senses beyond the traditional audiovisual scenes.
To tackle the aforementioned issues, we propose an architecture that harnesses semantic communication through Generative AI (GenAI) to facilitate immersive communication. Specifically, the code-generation capability of GenAI is employed to replicate the video captured by the UAV, thereby minimizing bandwidth consumption. This approach provides the drone teleoperator with a secondary stream that can be utilized in case the primary video stream is interrupted. Moreover, GenAI will also facilitate the generation of additional senses beyond the conventional audiovisual scenes, thus paving the way for the IoS.
IV-B Proposed Architecture for GenAI Enabled Immersive Communication

### v2

@@ Abstract
Over the past two decades, the Internet-of-Things (IoT) has become a transformative concept, and as we approach 2030, a new paradigm known as the Internet of Senses (IoS) is emerging. Unlike conventional Virtual Reality (VR), IoS seeks to provide multi-sensory experiences, acknowledging that in our physical reality, our perception extends far beyond just sight and sound; it encompasses a range of senses. This article explores the existing technologies driving immersive multi-sensory media, delving into their capabilities and potential applications. This exploration includes a comparative analysis between conventional immersive media streaming and a proposed use case that leverages semantic communication empowered by generative Artificial Intelligence (AI). The focal point of this analysis is the substantial reduction in bandwidth consumption by 99.93% in the proposed scheme. Through this comparison, we aim to underscore the practical applications of generative AI for immersive media. Concurrently addressing major challenges in this field, such as temporal synchronization of multiple media, ensuring high throughput, minimizing the End-to-End (E2E) latency, and robustness to low bandwidth while outlining future trajectories.
@@ I Introduction
The advent of the 5th generation (5G) mobile networks and recent advancements in computing technologies have redefined the concept of Internet from basic connectivity to a more advanced digital experience, transitioning from merely faster communication into an immersive interaction with the digital realm. This concept has been recently introduced under the umbrella of Metaverse and Digital Twins. It has opened up a wide range of applications including VR, Augmented Reality (AR), holoportation, and teleoperation, among others. Within this realm, four main underpinnings have been remarked as paradigms for linking the cyber and physical worlds, namely, connected intelligent machines, a digitized programmable world, connected sustainable world, and the IoS  [^1] https://www.ericsson.com/en/reports-and-papers/consumerlab/reports/10-hot-consumer-trends-2030. The IoS concept is set to revolutionize the digital interactions by creating a fully immersive environment that transcends traditional boundaries. By integrating sensory experiences such as sight, sound, touch, smell, and taste into the digital realm, this technology promises a more engaging cyber world, where virtual experiences are as rich and multi-dimensional as the physical world.
@@
Humans being experiencing the world through different senses, by perceiving sensory signals that are integrated or segregated in the brain. If these senses, especially haptic feedback, are accurately represented to be coherent with the real world, they can positively influence actions and behaviors, such as reaction time and detection [1]. Within this context, the IoS technology will allow individuals to experience a wide range of sensations remotely, revolutionizing various verticals, including industry, healthcare, networking, education, and tourism, to name a few. In order to reap the full potential of the IoS technology, numerous challenges need to be tackled to achieve a fully immersive multisensory experience. These challenges are pertinent to the temporal synchronization of multiple media, addressing motion sickness, ensuring high throughput, and minimizing the E2E latency. The collection of data from various sensor modalities, such as visual, audio, and haptic, plays a vital role in crafting a multisensory experience, in which this data can be synchronized at either the source or the destination (i.e., end devices or edge servers). The failure of virtual experiences to truly replicate our senses introduces confusion in human brains, leading to symptoms like nausea, dizziness, and migraines. To mitigate these drawbacks, it is crucial to enhance the realism of virtual sensations and reduce latency in VR/AR devices, thereby minimizing latency between different modalities and avoiding its mismatch [2]. Furthermore, for accurate control purposes over a distance of up to one mile and to prevent the occurrence of motion sickness, it is crucial to transmit the sensory information at extremely low E2E latency, ideally within 1-10 millisecond [3].
@@
With respect to the Key Performance Indicators for reliable communication of immersive media in IoS, it was demonstrated that future 6G networks should realize an E2E latency performance within the range of 1 ms for high-quality video streaming and haptic signals, with data rate requirements ranging from tens of Mbps to 1 Tbps and reliability performance of $10^{-7}$ [4]. In addition, while taste and smell signal requirements are less stringent than videos and haptics, it is essential to realize a perfect synchronization among signals from different senses to achieve the full potential of the IoS. Among various technologies, semantic communication emerges as a promising candidate for achieving ultra-low latency communication through communicating the meanings/semantics of messages instead of communicating the physical signal, yielding faster and bandwidth-efficient transmission.
@@
In this paper, we aim to set the scene for the integration of LLMs and the IoS technology, in which we develop a case study to demonstrate the benefits that can be obtained from exploiting the capabilities of LLMs in enhancing the latency performance of immersive media communication. In particular, we conceptualize the 360 video streaming from a Unmanned Aerial Vehicle (UAV) as a semantic communication task. Initially, we employ object detection and image-to-text captioning to extract semantic information (text) from the input 360 frame. Subsequently, this generated textual information is transmitted to the edge server. In the edge server, an LLM is utilized to produce WebXR code, facilitating the display of the corresponding image through Three-Dimensional (3D) virtual objects on the Head Mounted Device (HMD), and estimate Multi-Sensorial Media (Mulsemedia) sensors that actuate wearables to mimic the real environment’s thermal and haptic sensations. Lastly, the generated Mulsemedia and code are sent to the receiver, allowing for the rendering of the 3D virtual content on the HMD and direct actuation of haptic and thermal devices. The contributions of this paper are summarized as follows:
@@
  • Assess the quality of the generated 3D objects from our system compared to the captured 360 video images using reverse image-to-text, followed by text comparison through Bidirectional Encoder Representations from Transformers (BERT) model.
The remainder of this paper is organized as follows. Section II introduces the IoS and discusses its necessity. In Section III, an overview of the development of MLLMs and their applications to IoS is discussed. Section IV explores the state of the art in immersive media streaming. Section V presents a case study with a proposed testbed, which is implemented and analyzed. Section VI presents the experimental results. Finally, Section VIII highlights challenges and suggests directions for future research.
Figure 1: Key concepts of IoS
@@ II-A Immersive All-Sense Communication
Internet-of-Touch. Haptic sensation refers to the sense of touch, known as tactile sensation, and it enhances immersive multimedia by allowing individuals to feel physical sensations, such as interactions with objects and movements (kinesthetic sensation). In VR training or teleoperation, haptics replicate touch, which is crucial for tasks such as surgery. Achieving optimal haptic experiences requires addressing minimal response times and low latency in synchronization with other sensed media, such as audio and video. Haptic interfaces employ various technologies to deliver tactile sensations, ranging from simple vibration feedback to more complex systems providing force feedback, pressure sensitivity, or even localized temperature changes. Devices including haptic gloves, exoskeletons, or tactile feedback controllers enable users to touch, grasp, and interact with virtual objects in a natural and intuitive manner.
Internet-of-Taste. Gustatory perception involves the intricate process of detecting and interpreting flavors. While traditional VR primarily focuses on visual and auditory stimuli, incorporating taste into the virtual environment has the potential to enhance sensory engagement, leading to more realistic and immersive experiences. The technology underlying gustatory interfaces centers on the controlled stimulation of taste receptors. Various approaches are being explored, such as electrically stimulating taste buds [8] or delivering taste-related chemical compounds directly to the mouth.However, it is crucial to note that replicating the sense of taste is the most complex, as it closely depends on other sensations. Presently, the technology is still in the laboratory demonstration stage.
@@ III-A Advancement of Language Models
The progress in Natural Language Processing (NLP) research has led to the development of models such as GPT-2, BART  [^2] https://huggingface.co/docs/transformers/en/model_doc/bart, and BERT  [^3] https://huggingface.co/docs/transformers/en/model_doc/bert These models have sparked a new race to construct more efficient models with large-scale architectures, encompassing hundreds of billions of parameters. The most popular architecture is the decoder-only, including LLMs like GPT-3, Chinchilla and LaMDA. Following the release of open-source LLMs like OPT and BLOOM, more efficient open-source models have been recently introduced, such as Vicuna, Phi-1/2, LLaMa, FALCON, Mistral, and Mixtral444 https://huggingface.co/docs/transformers/model_doc/mixtral. This later follows the Mixture of Expects (MoE) architecture and training process initially proposed in MegaBlocks 555 https://huggingface.co/papers/2211.15841. Despite having fewer parameters, these models fine-tuned on high-quality datasets, have demonstrated compelling performance on various NLP tasks, surpassing their larger counterparts. Furthermore, instruction tuning the foundation models on high-quality instruction datasets enables versatile capabilities like chat and code source generation, etc. The LLM have also shown unexpected capabilities of learning from the context (i.e., prompts), referred to as In-Context Learning (ICL).
@@ III-B Multimodal large language models
In this era of MLLMs, GPT-4 has demonstrated remarkable performance in vision-language tasks encompassing comprehension and generation. Nevertheless, in addition to its intricate nature, the technical details of GPT-4 remain undisclosed, and the source code is not publicly available, impeding direct modifications and enhancements.To address these challenges, the MiniGPT-4 model was proposed. This model combines a vision encoder (ViT-G/14 and Q-Former) with the Vicuna LLM, utilizing only one projection layer to align visual features with the language model while keeping all other vision and language components frozen. The model is first trained on image-language datasets, then finetuned on high-quality image description pairs (3,500) to improve the naturalness of the generated language and its usability. The TinyGPT-V vision model, introduced by Yuan et al. [12], addresses computational complexity, necessitating only a 24GB GPU for training and an 8GB GPU for inference. The architecture of TinyGPT-V closely resembles that of MiniGPT-4, incorporating a novel linear projection layer designed to align visual features with the Phi-2 language model, which boasts only 2.7 billion parameters. The TinyGPT-V model undergoes a sophisticated training and fine-tuning process in four stages, where both the weights of the linear projection layers and the normalization layers of the language model are updated. As the process progresses, instruction datasets are incorporated in the third stage, and multi-task learning is employed during the fourth stage.
Figure 2: Architecture of a conventional video streaming system
The second step in developing LLMs is fine-tuning the model on instruction datasets to teach models to better understand human intentions and generate accurate responses. The InstructBLIP 666 https://huggingface.co/docs/transformers/model_doc/instructblip is built through instruct tuning of the pre-trained BLIP-2 model on 26 instruction datasets grouped into 11 tasks. During the instruction tuning process, the LLM and the image encoder are maintained frozen, while only the Q-former undergoes fine-tuning. Furthermore, the instructions are input to both the frozen LLM and the Q-Former. Notably, InstructBLIP exhibits exceptional performance across various vision-language tasks, showcasing remarkable generalization capabilities on unseen data. Moreover, when employed as the model initialization for individual downstream tasks, InstructBLIP models achieve state-of-the-art fine-tuning performance. InstructionGPT-4 777 https://huggingface.co/datasets/WaltonFuture/InstructionGPT-4 is a vision model fine-tuned on a small dataset comprising only 200 examples, which represents approximately 6% of the instruction-following data used in the alignment dataset for MiniGPT-4. The study highlights that fine-tuning the vision model on a high-quality instruction dataset enables the generation of superior output compared to MiniGPT-4.
@@
IV State of the Art of conventional and semantic immersive media streaming Methods
TABLE I: Comparison of Different Papers on Streaming Technologies
Category Paper Type Media Type Streaming Protocol Encoder / Decoder Design Objective Conventional Live Streaming Lo et al. (2018) [13] VoD 360 video ERP frames DASH HEVC/H.265 codec Bandwidth and latency Taleb et al. (2022) [14] Live 360 video ERP frames WebRTC AVC/H.264 codec Ultra low latency Chen et al. (2021) [15] Live 360 video ERP tiles DASH AVC/H.264 codec Bandwidth $>$ 50% traditional streaming Yi et al. (2020) [16] Live 360 video ERP frames RTMP over HTTP-FLV AVC/H.264 codec Latency $<$ 4s for 1440p resolution Park et al. (2023) [17] Live Spherical to 360 video ERP RTMP over HLS AVC/H.264 codec Super resolution and bandwidth saving Gao et al. (2024) [18] Live 360 video ERP tiles RTMP over LL-DASH AVC/H.264 codec Scalability De Fré et al. (2024) [19] Live Head position + 3D video WebRTC Draco codec 360ms latency for 1080p video at 3Mb/s Usón et al. (2024) [20] Live Volumetric video WebRTC V-PCC codec Optimal latency at 70Mb/s bandwidth Semantic Streaming Xia et al. (2023) [21] VoD 360 video tiles / CNN Encoder / Decoder Reduce latency with reliable transmission Ahn et al. (2024) [22] VoD Video over text semantics / GPT-4 Encoder / DALLE-2 Decoder Video content creation Chen et al. (2024) [23] VoD Text, audio, image, haptics / GNN Encoder/3D generative reconstruction network Decoder 3D object construction Du et al. (2023) [24] VoD 3D objects through text / CNN (YOLOv7) Encoder / Database Decoder Optimize transmission power Ours Live Text, video frames, temperature, and haptics HTTP for semantics, MQTT for sensorial data and generated code CNN + RNN + GPT-3.5 Encoder GPT-4 Decoder Optimize bandwidth consumption
The latest implementations and research on live immersive media streaming typically adhere to the conventional pipeline illustrated in Fig. 2. This pipeline involves capturing a scene using either a 360 camera or multiple cameras, followed by stitching the frames and encoding. The encoding can occur either at the camera itself or on a separate processing unit. Subsequently, the frames are projected into an Equirectangular Projected (ERP) format or cube map and encoded using traditional video standards such as AVC/H.264 or HEVC/H.265. Due to the resource limitations of 360 cameras, the encoded stream is usually transmitted to a remote media server using Real-Time Messaging Protocol (RTMP) or Real-Time Streaming Protocol (RTSP). The media server may then re-encode the video before transmitting it to the end-user via Dynamic Adaptive Streaming over HTTP (DASH), Web Real-Time Communication (WebRTC), or another media streaming protocol. Previous studies have shown WebRTC to be particularly effective due to its ultra-low latency and adaptive bitrate capabilities [25]. For Video on Demand (VoD) services, the primary distinction lies in the storage of the video on a cloud server instead of real-time transmission. For point cloud video, alternative codecs such as Google Draco or Video-based Point Cloud Compression (V-PCC) are employed..
Recent research has explored the use of AI as a compressor for semantic communication, transmitting only essential knowledge and information for scene reconstruction at the receiver. This approach holds promise for reducing redundant data and conserving bandwidth, proving particularly advantageous in high-mobility, frequent-handover scenarios like UAV communication and control. Table I summarizes recent studies comparing traditional streaming pipelines with semantic communication approaches in terms of protocols, codecs, and design. While some traditional techniques incorporate Field of View (FoV) prediction and tile encoding for bandwidth optimization, they still operate in the megabits per second range. This limitation can result in video feed loss in severely bandwidth-constrained environments, a challenge not yet addressed by existing methods. Furthermore, current semantic communication-based solutions often remain confined to simulations and are not tailored for real-time applications. Our proposed architecture, to our knowledge, is the first Generative AI (GenAI)-based encoder/decoder for immersive multimedia streaming in a real-time, ultra-low latency application like UAV control. We have chosen the solution in [14] as a benchmark because it represents one of the optimal traditional pipelines based on WebRTC, achieving ultra-low E2E latency $\leq$ 600ms) for immersive streaming in UAV control scenarios.
V Case Study
V-A Use case description
@@
To comprehend the challenges at hand and explore potential solutions, let us immerse ourselves in the following scenario. John, a surveillance teleoperator, is tasked with remotely piloting a drone through a dense forest using a First-Person View (FPV) system over a Beyond 5G (B5G) network. The task of navigating this complex environment through FPV poses significant difficulties due to two primary factors:
  • Limited Bandwidth: The forest environment inherently restricts bandwidth, leading to a degraded video stream in John’s FPV system. This degradation impairs his ability to effectively control the drone, potentially resulting in hazardous situations.
  • Limited Sensory Input: The drone’s 360-degree camera, while providing visual and auditory feedback through the FPV system, fails to fully capture the rich sensory context of the UAV’s surroundings. Achieving a truly immersive and comprehensive understanding of the drone’s environment would require additional sensory inputs beyond the traditional visual and auditory sensors.
To address these challenges, we propose an architecture that leverages GenAI for semantic communication. This approach aims to:
  • Reduce Bandwidth Consumption: GenAI’s code-generation capabilities can be used to replicate the drone’s video feed, minimizing bandwidth usage. This provides John with a secondary video stream, ensuring continuous operation even if the primary stream is interrupted.
  • Enhance Sensory Immersion: GenAI can generate additional sensory information beyond the traditional visual and audio streams, paving the way for an IoS experience. This will allow John to perceive the environment more comprehensively, improving his ability to control the drone safely and effectively.
By implementing this architecture, we can create a more immersive and reliable remote drone operation system, enabling teleoperators to navigate challenging environments with greater confidence and precision.
Furthermore, DT based on 3D simulated environments have received a lot of interest from researchers, specifically for UAV teleoperation. [26] proposes a DT framework for UAV monitoring and autonomy in which the UAV executes missions only after successful simulation of the UAV in the DT. [27] present a framework for UAV control through VR comprising a DT UAV equipped with virtual sensors that override user commands if obstacles in the DT environment are detected nearby, thus providing reliable teleportation. However, all of those DT-based solutions rely on static 3D maps, which tend to evolve over time with the incorporation of temporary objects, thus making the static DT unreliable. Therefore, we solve the latter issue by leveraging our proposed architecture. We enable UAVs to capture detailed environmental data, specifically of temporary elements in the environment that are difficult to find in any 3D database. Thus, we are able to inject the temporary objects generated by GenAI, as proposed in our framework, into the DT, and save bandwidth by streaming only the changing elements of the environment.
V-B Proposed Architecture for GenAI Enabled Immersive Communication